# KNighter: Transforming Static Analysis with LLM-Synthesized Checkers

### v1

@@ 3.1.2 Plan Synthesis
After identifying the bug pattern, KNighter synthesizes a high-level plan for implementing the static analyzer. This planning phase is critical for two reasons: it provides structured guidance for LLMs during implementation, thereby preventing confusion or ineffective execution, and it facilitates debugging of the entire pipeline by making the LLMs’ reasoning process transparent and traceable. Our ablation study in § 5.4.2 demonstrates that having a plan leads to better performance than not having one, consistent with findings from other domain applications [27].
Figure 5: Prompt template for plan synthesis.
@@
To synthesize the implementation plan for the checker, we have designed an LLM-based agent whose prompt template is shown in Figure 5. This agent takes the previously summarized bug pattern as input. Additionally, we maintain a curated database of utility functions for checker implementation that can be easily extended. By including the signatures and brief descriptions of these utility functions in the prompt, we enable LLMs to leverage them effectively during the planning process, simplifying the overall task.
3.1.3 Analyzer Implementation and Syntax Repair
@@
After identifying the bug pattern and making the plan, we leverage an LLM-based agent to implement the corresponding checker, as illustrated in Figure 6. To maximize implementation accuracy, we provide the agent with comprehensive inputs: the distilled bug pattern and a structured implementation plan. We also provide a pre-defined checker template that standardizes the implementation structure and reduces potential errors. Moreover, we provide a list of utility functions that could help with the implementation.
@@
To handle potential compilation errors in the generated checkers, we employ a dedicated debugging agent. Inspired by existing work on program repair [32], this agent automatically processes compiler error messages and applies necessary fixes, effectively addressing syntax errors that may arise from LLM hallucinations. This automated debugging pipeline ensures that the final checkers are both syntactically correct and compilable.
Figure 6: Prompt template for analyzer implementation.
3.1.4 Validation
To validate our checkers semantically and mitigate hallucinations by LLMs, we evaluate them against both the buggy (pre-patch) and patched versions of the Linux kernel. This dual-version analysis verifies that the checker can both detect original bugs and recognize their fixes. For efficiency, we restrict the evaluation to only the files modified by the patch and their dependencies, rather than scanning the entire kernel codebase. This focused approach enables rapid validation while maintaining effectiveness. We assess checker validity by comparing the number of reported issues between versions, acknowledging that while similar bugs might persist in the patched version, an effective checker should yield a meaningful difference in detection results. Detailed implementation specifics are presented in § 4.
3.2 Checker Refinement
After synthesizing a valid checker, we scan the entire Linux codebase. While effective at identifying potential issues, these checkers may be overly conservative, flagging correct code as problematic and generating false positives. To address this limitation, we develop an iterative refinement process powered by LLMs. The process evaluates each bug report’s validity and uses false positive cases to refine the checker. However, this refinement process faces two significant challenges:
  • Bug reports often contain extensive context and implementation details, making them difficult to process efficiently.
  • Debugging and refining checkers based on false positive reports requires complex analysis and careful modification.
To address these challenges, we first distill bug reports to their essential elements. We extract only the "relevant lines" identified by the LLVM report and the trace path for each potential issue—preserving critical diagnostic information while eliminating extraneous context. We then deploy a specialized LLM-based triage agent to classify each report as either a true or false positive. Note, our classification criteria focus specifically on pattern alignment: we evaluate whether each reported issue exhibits the same bug pattern as the original input patch, rather than conducting a general correctness assessment. For reports classified as false positives, a dedicated refinement agent analyzes the specific case and modifies the checker accordingly, enhancing precision while preserving the checker’s ability to detect the original vulnerability pattern.
@@ 3.2 Checker Refinement
4 Implementation
Input commit collection. To collect patch commits, we implemented a systematic classification and selection process. First, we established 10 distinct bug categories. We then used relevant keywords to identify potentially related commits. A commit was included in our dataset only when two authors independently agreed on its categorization. For each bug type, we initially examined the first 20 commits that matched our search criteria. We continued reviewing commits beyond the initial 20 if we hadn’t yet collected 5 qualifying commits for a given category. Our goal was to gather a minimum of 5 commits per bug type whenever possible. Table 1 presents our categorization of 10 bug types and their corresponding patch commit counts.
Table 1: Distribution of patch commits across 10 bug categories and the validity status of their synthesized checkers. “NPD” denotes “Null-Pointer-Dereference” and “UBI” indicates “Use-Before-Initialization”.
@@ 4 Implementation
Few-shot examples. We prepared three end-to-end examples for LLMs’ in-context learning. These three are patch commit 3027e7b15b02 (Null-Pointer-Dereference), 3948abaa4e2b (Use-Before-Initialization), and 4575962aeed6 (Double-Free). The design and implementation of the checker for these three commits required approximately 40 person-hours.
Utility functions. While implementing checkers for the example commits, we identified several commonly needed utility functions. We implemented 9 utility functions, including getMemRegionFromExpr, to support checker development. These utility functions are designed to be simple to implement and extend.
Valid checkers. To evaluate checker validity, we verify that it can both detect the original bug and recognize its fix. We first identify buggy objects by examining the modified files in the diff patch. Next, we check out the repository to the buggy commit (immediately preceding the patch) and scan these objects to count the number of bug reports ($N_{buggy}$). We then scan the same objects after applying the patch commit to obtain the number of remaining bug reports ($N_{patched}$). A checker is considered valid if $N_{buggy}>N_{patched}$ and $N_{patched}<T_{valid}$, where $T_{valid}$ is a threshold value, which is 50 by default.
@@ 5 Evaluation
    Can KNighter synthesize high-quality static analyzers?
@@
    How do the static analyzers synthesized by KNighter compare with human-written ones?
@@
  • Runtime Errors: Checkers that compile successfully but crash during execution (e.g., "The analyzer encountered problems on source files").
@@
Hardware and software. All our experiments are run on a workstation with 64 cores, 256 GB RAM, and 4 Nvidia A6000 GPUs, operating on Ubuntu 20.04.5 LTS. We use O3-mini as our default LLM backend. By default, when scanning the entire codebase, we use -j32. We evaluated using Linux v6.13, and for bug finding, we examined versions from v6.9 to v6.13. The Linux configuration used is allyesconfig.
5.1 RQ1: Synthesized Checkers
(a) The bug in the Use-Before-Initialization patch.
(b) A false positive reported by KNighter.
Figure 7: Reported false positive for a UBI commit.
Synthesis. From the commits collected in Table 1, we successfully generated valid checkers for 39 commits. Furthermore, KNighter successfully generated valid checkers across diverse bug types beyond those in our few-shot examples, demonstrating the generalizability of our approach.
The complete synthesis process required 15.9 hours, consuming 8.2 million input tokens and producing 1.2 million output tokens. For commits yielding valid checkers, KNighter required an average of 2.4 attempts per checker (maximum: 8 attempts). The generated valid checkers averaged 125.7 lines of code.
@@ 5.1 RQ1: Synthesized Checkers
Refinement. After scanning the entire kernel codebase with these 39 valid checkers, 26 of them were labeled “plausible” directly. Our refinement pipeline was applied to the remaining 13 valid checkers, successfully refining 11 of them. In total, 19 refinement steps were completed successfully. This demonstrates the effectiveness of our refinement pipeline, which successfully refined 84.6% of the valid checkers that were not initially labeled as “plausible”.
@@
Here are some example checkers with poor performance (high false positive rate). For the commit 90ca6956d383 (“ice: Fix freeing uninitialized pointers”) shown in Figure 7(a), the issue occurs when the pointer pcaps is not set to NULL, and an early return or error path is taken before it is assigned a valid allocation, causing the cleanup logic to attempt to free an uninitialized (or garbage) pointer. However, to prove such a bug pattern exists, there must be an early exit-path where the pointer is not assigned any value. Thus, for the bug report shown in Figure 7(b), although the cert pointer is not initialized, it will be directly assigned a value and there is no path where it remains unassigned. Therefore, this is not a bug. Our synthesized checker failed to consider this constraint, and our triage agent also overlooked this issue.
Invalid checkers. During the checker synthesis process (up to 10 attempts for each commit), we recorded in total 273 failed attempts across all 61 commits. These invalid checkers failed for three primary reasons: 65 encountered compilation errors, 1 exhibited runtime errors, and the remaining 207 contained semantic issues that prevented proper bug identification.
Of the 207 checkers with semantical issues, 34 incorrectly labeled both buggy and patched code as potential bugs, while the remaining 173 incorrectly labeled both versions as not containing bugs. This demonstrates that correctly identifying buggy code remains a significant challenge for KNighter.
Interestingly, the 173 checkers that fail to identify the specific bug in their input patch can still provide value. When applied to the broader kernel codebase, these checkers may successfully detect bugs with similar patterns in other contexts. This apparent paradox occurs because a checker’s failure to detect its training bug might be due to specific edge cases or context-dependent factors rather than fundamental flaws in its detection logic. For instance, a checker might not detect a null pointer dereference in its training example due to complex control flow, but still correctly identify simpler instances of the same pattern elsewhere. Moreover, these checkers often have lower false positive rates compared to those that incorrectly flag both buggy and patched code as problematic, making them even more practical for real-world bug detection.
Our approach shows limitations in detecting buffer overflow and use-after-free vulnerabilities. A key challenge is that CSA has inherent difficulty in performing precise value analysis, particularly in determining buffer bounds at compile time. This limitation makes it challenging to statically reason about potential overflow conditions that depend on runtime values.
Table 2: Detected bugs by KNighter.
@@
KNighter 70 56 41 15 11
(a) Number of bugs in each subsystem.
@@
Figure 8: The details of newly detected bugs by KNighter.
@@
Figure 9: Example vulnerabilities detected by KNighter.
5.2 RQ2: Detected Bugs
To date, the static analyzers generated by KNighter have detected 70 new bugs, with 56 confirmed and 41 already fixed (see Table 2). Notably, 11 of these bugs have been assigned CVE numbers. We identified more bugs than those reported as true positives in RQ-1 because our analysis spanned a broader range of Linux kernel versions.
Figure 8 shows the details of newly detected bugs, including their subsystem locations and bug types. We found that most bugs are from the drivers subsystem (45/70), which aligns with the fact that driver code constitutes the largest portion of the kernel [6]. Following drivers, we also found 12 and 6 bugs in the sound and net subsystems respectively.
As shown in Figure 8(b), KNighter demonstrates capability in detecting a diverse range of bug types. Null-Pointer-Dereference issues constituted the largest category (30/70 bugs), followed by Integer-Overflow and Misuse vulnerabilities. This distribution likely reflects the pattern-based nature of Null-Pointer-Dereference and Misuse bugs, which makes their checkers more amenable to LLM-based implementation. Integer-Overflow bugs, typically involving variable type mismatches, benefit from the Clang toolchain’s strong type analysis capabilities. More complex issues such as Use-After-Free and Concurrency bugs proved more challenging to detect at compile time, particularly those involving multi-threaded execution contexts. Overall, KNighter-synthesized checkers are general to detect various types of bugs.
Here are example vulnerabilities detected by KNighter.
@@ 5.2 RQ2: Detected Bugs
CVE-2024-50259. Figure 9(c) shows an input patch addressing a buffer overflow vulnerability. The patch mitigates the risk by limiting the number of bytes copied via copy_from_user to sizeof(mybuf) - 1, thereby preserving space for a trailing zero. This trailing zero is essential for subsequent string operations, such as sscanf, to function correctly. Leveraging this patch as a reference, the checker generated by KNighter identified a similar bug in nsim_nexthop_bucket_activity_write, as shown in Figure 9(d). In this case, the omission of appending a trailing zero after copying data from userspace could lead to improper string handling and potential overflow issues. This detected issue was subsequently fixed by adding the trailing zero and was assigned CVE-2024-50259.
5.3 RQ3: Baseline Comparison

### v2

@@ 3.1.2. Plan Synthesis
Once the bug pattern is identified, KNighter generates a high-level plan for implementing the static analyzer. This plan serves two critical purposes: first, it provides structured guidance to the LLMs during implementation, preventing confusion and promoting effective execution. Second, it facilitates debugging of the entire pipeline by making the LLMs’ reasoning process transparent and traceable. Our ablation study in § 5.4.2 confirms the value of this plan synthesis, demonstrating improved performance consistent with findings in other domains (sun2023pearl, ).
For instance, synthesizing a checker for the unchecked devm_kzalloc return value pattern (illustrated in Figure 2(c)) might generate a plan with key steps such as: (1) Using program state to track memory regions from devm_kzalloc, (2) monitoring conditional branches (checkBranchCondition) to mark regions as checked if a null check occurs, and (3) detecting uses (checkLocation) of unchecked regions, potentially signaling a bug. This high-level structure guides the subsequent implementation phase.
@@
To synthesize the implementation plan for the checker, we have designed an LLM-based agent whose prompt template is shown in Figure 5(b). This agent takes the previously summarized bug pattern as input. Additionally, we maintain a curated database of utility functions for checker implementation that can be easily extended. By including the signatures and brief descriptions of these utility functions in the prompt, we enable LLMs to leverage them effectively during the planning process, simplifying the overall task.
(a) Prompt template for bug pattern analysis
(b) Prompt template for plan synthesis.
Figure 5. Prompt templates used by KNighter.
3.1.3. Implementation and Syntax Repair
@@
After identifying the bug pattern and making the plan, we leverage an LLM-based agent to implement the corresponding checker. To maximize implementation accuracy, we provide the agent with comprehensive inputs: the distilled bug pattern and a structured implementation plan. We also provide a pre-defined checker template, as shown in Figure 6, which standardizes the implementation structure and reduces potential errors. Moreover, we provide a list of utility functions that could help with the implementation.
@@
The synthesized checkers could have compilation errors, e.g., using the wrong static analysis API or incorrect variable types. To handle the potential compilation errors, we employ a dedicated debugging agent. Inspired by existing work on program repair (xia2024automated, ), this agent automatically processes compiler error messages and applies necessary fixes, effectively addressing syntax errors that may arise from LLM hallucinations. This automated debugging pipeline ensures that the final checkers are both syntactically correct and compilable.
Figure 6. Pre-defined checker template for CSA.
3.1.4. Validation
To semantically validate our checkers and mitigate potential inaccuracies from LLMs, we evaluate them against both the buggy (pre-patch) and patched versions of the relevant code (e.g., Linux kernel files). This differential analysis verifies that a checker correctly identifies the target bug in the original code and confirms its absence after the patch. For efficiency, we scope this validation to only the files modified by the patch and their dependencies, rather than the entire codebase. A checker is considered valid if it flags the bug in the pre-patch version and shows a corresponding reduction or elimination of that specific warning in the patched version. More details are in § 4.
Figure 7. A report labeled as FP by our triage agent.
3.2. Checker Refinement
Following synthesis, each valid checker is used to scan the entire system. However, its initial validation doesn’t prevent potential false positives when analyzing the broader codebase, where correct code might be flagged erroneously. To mitigate this, we implement an iterative refinement procedure driven by LLMs. This involves evaluating the generated bug reports and feeding identified false positives back to refine the checker.
However, automating this refinement faces hurdles. First, bug reports are often verbose, containing extensive context that is difficult to process efficiently. Second, debugging the checker logic and modifying it correctly based on false positives requires complex analysis.
Our refinement pipeline addresses these challenges methodically. First, to manage report complexity, we distill generated bug reports to their essential components—primarily the “relevant lines” highlighted by the static analyzer (e.g., CSA (clang-static-analysis, )) and the corresponding trace path—stripping extraneous context while preserving critical diagnostics. Second, to navigate the complexity of analysis and modification, we employ specialized LLM-based agents. A triage agent classifies each distilled report, focusing strictly on pattern alignment (i.e., does the report match the original input bug’s pattern?) rather than general code correctness.
If the triage agent identifies a report as a false positive, as exemplified in Figure 7, a dedicated refinement agent takes over. In the case shown, the initial checker (derived from the patch in Figure 2(a)) flagged the use of pmx->pfc because its logic failed to recognize if (unlikely(!pmx)) as a valid null check, perhaps confused by the unlikely() macro. The triage agent correctly interprets the check semantically and flags the report as FP. The refinement agent then uses this information to adjust the checker’s logic, specifically enhancing its ability to handle constructs like unlikely(), thereby preventing this type of false positive in subsequent scans while ensuring it can still detect the original vulnerability.
@@ 3.2. Checker Refinement
4. Implementation
Input commit collection. To collect patch commits for rigorous evaluation, we implemented a systematic classification and selection process. First, we established 10 distinct bug categories. We then used relevant keywords to identify potentially related commits. A commit was included in our dataset only when two authors independently agreed on its categorization. For each bug type, we initially examined the first 20 commits that matched our search criteria. We continued reviewing commits beyond the initial 20 if we hadn’t yet collected 5 qualifying commits for a given category. Our goal was to gather a minimum of 5 commits per bug type whenever possible. Table 1 presents our categorization of 10 bug types and their corresponding patch commit counts. These manually collected and labeled commits served as our benchmark dataset for rigorous evaluation.
Table 1. Distribution of patch commits across 10 bug categories and the validity status of their synthesized checkers. “NPD” denotes “Null-Pointer-Dereference” and “UBI” indicates “Use-Before-Initialization”.
@@ 4. Implementation
Few-shot examples. We prepared three end-to-end examples for in-context learning. These three are patch commit 3027e7b15b02 (Null-Pointer-Dereference), 3948abaa4e2b (Use-Before-Initialization), and 4575962aeed6 (Double-Free). The design and implementation of the checker for these three commits required approximately 40 person-hours. This was a one-time effort, yielding reusable examples. We also explore the use of real-world, off-the-shelf examples (§ 5.4.2).
Utility functions. While implementing example checkers, we identified several common helper operations. We implemented 9 such utility functions (e.g., getMemRegionFromExpr) to encapsulate low-level Clang Static Analyzer tasks, simplifying checker development, particularly for LLM synthesis. These utilities were designed for simplicity and extensibility.
Valid checkers. To evaluate checker validity, we verify that it can both detect the original bug and recognize its fix. We first identify buggy objects by examining the modified files in the diff patch. Next, we check out the repository to the buggy commit (immediately preceding the patch) and scan these objects to count the number of bug reports ($N_{buggy}$). We then scan these objects after applying the patch commit to obtain the number of remaining bug reports ($N_{patched}$). A checker is considered valid if $N_{buggy}>N_{patched}$ and $N_{patched}<T_{valid}$, where $T_{valid}$ is a threshold value (50 by default).
@@ 5. Evaluation
    Can KNighter generate high-quality checkers?
@@
    Are the capabilities of KNighter orthogonal to the human-written checkers?
@@
  • Runtime Errors: Checkers that compile successfully but crash during execution (e.g., ”The analyzer encountered problems on source files”).
@@
Static Analysis Capabilities. We further examine the static analysis capabilities employed by checkers, including path sensitivity, region sensitivity, and advanced state tracking.
@@
Hardware and software. All our experiments are run on a workstation with 64 cores, 256 GB RAM, and 4 Nvidia A6000 GPUs, operating on Ubuntu 20.04.5 LTS. We use O3-mini as our default LLM backend. By default, when scanning the entire codebase, we use -j32. We evaluated using Linux v6.13, and for bug finding, we examined versions from v6.9 to v6.15. The Linux configuration used is allyesconfig.
5.1. RQ1: Synthesized Checkers
(a) Bug in Use-Before-Initialization patch.
(b) False positive report for UBI.
Figure 8. Examples of false positives by KNighter.
We evaluate KNighter on the 61 commits listed in Table 1 to show that it can synthesize high-quality checkers across various bug types, beyond those in our few-shot examples.
5.1.1. Checker Synthesis
In total, valid checkers were generated for 39 commits. The 39 synthesized checkers average 125.7 lines of code and exhibit diverse static analysis capabilities. Specifically, 37 checkers are path-sensitive, 13 incorporate region sensitivity, and 16 employ advanced state tracking; in contrast, only 2 leverage AST travelers. This suggests that KNighter can generate complex analysis logic, not just simple pattern matching.
Cost. The full synthesis process required 15.9 hours, during which 8.2 million input tokens were processed and 1.2 million output tokens produced, resulting in an approximate cost of \$0.24 per commit using O3-mini. For commits that ultimately yielded valid checkers, an average of 2.4 synthesis attempts was necessary (with a maximum of 8 attempts observed).
Failure analysis. We now break down the failures from two perspectives: the underlying failure root causes and the observed failure symptoms during the synthesis process.
(i) Failure root causes. Among the 61 commits processed, 22 did not result in any valid checker. Our investigation into these failures indicates that:
  • 2 (9%) commits failed due to an inaccurate bug pattern,
  • 7 (32%) failed owing to an inaccurate plan, and
  • 13 (59%) were caused by inaccurate implementation.
For the implementation-related failures, a common issue was that compiler optimizations inlined certain function calls (e.g., strcp and memset), which prevented the checker from properly intercepting these calls.
Our approach exhibits limitations in handling buffer overflow and use-after-free commits. We believe these challenges stem from two main factors: static analysis inherently struggles with precise value determination—especially when establishing buffer bounds at compile time—and it also faces significant hurdles in analyzing multi-threaded code, for instance, when assessing the proper use of locks.
(ii) Failure symptoms. During synthesis (allowing up to 10 attempts per commit), 273 failed attempts were recorded from all 61 commits. The failures can be categorized as:
  • 65 attempts (23.8%) resulted in compilation errors,
  • 1 attempt (0.4%) led to a runtime error, and
  • 207 attempts (75.8%) suffered from semantic issues that obstructed proper bug identification.
Of the 207 semantic failures, 34 checkers erroneously flagged both buggy and patched code as potentially problematic, while the remaining 173 misclassified both versions as bug-free. This outcome underscores the challenge of accurately distinguishing buggy code.
Interestingly, even the 173 checkers that did not recognize the specific bug in their input patch can still be valuable. When deployed across the large system, these checkers may successfully detect bugs with similar patterns in other contexts. This apparent paradox likely arises because the failure to detect the training bug is sometimes due to edge cases or context-specific complexities rather than inherent deficiencies in the checkers’ detection logic. Moreover, these checkers generally exhibit lower false positive rates compared to those that incorrectly flag both buggy and patched code, enhancing their practical utility for bug detection.
5.1.2. Checker Refinement.
@@ 5.1.1. Checker Synthesis
After scanning the entire kernel codebase with these 39 valid checkers, 26 of them were labeled “plausible” directly. Our refinement pipeline was applied to the remaining 13 valid checkers, successfully refining 11 of them. In total, 19 refinement steps were completed successfully. This demonstrates the effectiveness of our refinement pipeline, which successfully refined 84.6% of the valid checkers that were not “plausible” initially.
@@
Our manual analysis of the 29 false positives revealed three recurring patterns leading to incorrect reports:
  • Inaccurate bug pattern: In 5 cases, although the checker correctly identified the original bug/patch scenario, the inferred bug pattern lacked the necessary precision for reliable detection across different contexts.
  • Incorrect pattern matching: For 6 reports, the checker correctly identified the bug pattern but applied it too broadly, flagging code segments that did not meet the specific constraints intended by the pattern.
  • Trigger condition mismanagement: The most common issue (18 reports) involved checkers where both the pattern and matching were correct, but the checker failed to manage trigger or state conditions properly (e.g., failing to recognize a pointer had already been validated before use).
Case study: A high false positive rate checker. In commit 90ca6956d383 (“ice: Fix freeing uninitialized pointers”, see Figure 8(a)), the issue stems from the pointer pcaps not being initialized to NULL. If an early return or error path occurs before the pointer is allocated, the cleanup routine may inadvertently attempt to free an uninitialized (or garbage) pointer. In such cases, the checker should account for the possibility of an early exit leaving the pointer unset. In contrast, the bug report in Figure 8(b) highlights a scenario where, although the cert pointer starts uninitialized, it is immediately assigned a valid value along every execution path, ensuring it is never left in an unassigned state. Thus, despite the initial uninitialized state, the code does not constitute a bug. Our synthesized checker did not incorporate these nuanced constraints, and the triage agent likewise failed to recognize the critical differences, ultimately leading to a false positive.
@@
KNighter 92 77 57 15 16
(a) Bug summary statistics detected by KNighter.
@@
(c) Number of bugs in each subsystem.
(d) Number of bugs with different lifetimes.
Figure 9. The details of newly detected bugs by KNighter. Subfigure (9(a)) summarizes overall status, while (9(b)), (9(c)), and (9(d)) show breakdowns by type, subsystem, and lifetime.
5.2. RQ2: Detected Bugs
5.2.1. Overall.
To date, static analyzers synthesized by KNighter have identified 92 new bugs in the Linux kernel. As summarized in Figure 9(a), developer confirmation has been received for 77 of these bugs. Among those confirmed, 57 have already been fixed. The remaining 15 bugs are currently awaiting developer review (calculated as Total - Confirmed). Notably, 16 of the discovered bugs have been assigned CVE numbers, showing their practical security impact.
@@ 5.2. RQ2: Detected Bugs
Figure 10. Example vulnerabilities detected by KNighter.
The checkers for bug detection originate from two sources: (i) the initial 61 manually collected commits used for evaluation in § 3.1 across diverse bug types (shown in Table 1), and (ii) an additional set of 100 commits automatically collected using keywords related to Null-Pointer-Dereference to further explore that specific bug class. Figure 9(b) and Figure 9(c) show the bug distribution from these sources—with the light blue for bugs from the initial evaluation set and the light purple for those from the auto-collected set.
Bug types. Analysis of checkers from the manually collected commits reveals that KNighter can find a wide range of bug types (in Figure 9(b)). Among these, Null-Pointer-Dereference bugs are the most prevalent, showing the KNighter’s strength in this area. In response, we expanded our effort by automatically collecting commits related to Null-Pointer-Dereference, which yielded an additional 30 bugs.
Bug location. As shown in Figure 9(c), the detected bugs span various Linux kernel subsystems. The majority appear in the drivers subsystem (62 out of 92), reflecting its large footprint in the kernel (bursey2024syzretrospector, ). Additionally, 14 and 7 bugs were identified in the sound and net subsystems, respectively. Notably, 2 bugs were found in the samples directory—an area that provides example usage for kernel developers and where correctness is especially critical.
Bug lifetime. Figure 9(d) illustrates the distribution of bug lifetimes. Notably, the average bug lifetime is 4.3 years, and 26 bugs had existed for over five years before detection. This indicates that the vulnerabilities uncovered by KNighter are difficult to detect and remain latent for extended periods, underscoring the effectiveness of our approach.
5.2.2. Case Study.
Here are examples of vulnerabilities detected by KNighter.
@@
CVE-2024-50259. Figure 10(c) shows an input patch fixing a buffer overflow vulnerability. The patch mitigates the risk by limiting the number of bytes copied via copy_from_user to sizeof(mybuf) - 1, thereby preserving space for a trailing zero. This trailing zero is essential for subsequent string operations, such as sscanf, to function correctly. Taking this patch as input, the checker generated by KNighter identified a similar bug in nsim_nexthop_bucket_activity_write, as shown in Figure 10(d). In this case, the omission of appending a trailing zero after copying data from userspace could lead to improper string handling and potential overflow issues. This detected issue was subsequently fixed by adding the trailing zero and was assigned CVE-2024-50259.
5.3. RQ3: Orthogonality with Smatch